OpenStudy (anonymous):

5x-4y=-18 what are the 3 points to graph

OpenStudy (anonymous):

infinitely many points start by picking 3 different x values

OpenStudy (anonymous):

I dont know how to do any of it thats why i asked.

OpenStudy (anonymous):

You could solve it for y. Then plug in 3 values a suggestion would be x = 0 x = 1 x = 2

OpenStudy (anonymous):

Let x = 0. Plug in 0 for x and solve for y. After this, let x equal any other two numbers and solve for y. Here is the first one \[5(0)-4y=-18; -4y = -18\] 4y = 18 y = 18/4 y = 9/2 So a point on the graph would be (0,9/2)